OpenStudy (anonymous):

Choose the procedure you would use to solve the equation 5x + 6 = 41 for x.

OpenStudy (dmezzullo):

what math class is this?

OpenStudy (stamp):

Two-step algebra equation. @dmezzullo likely Algebra I

OpenStudy (stamp):

@dmezzullo or Pre-Algebra

OpenStudy (dmezzullo):

Ok ik how to do dis den

OpenStudy (anonymous):

5X=41-6 5X=35 X=7

OpenStudy (anonymous):

Well to get x alone you must isolate it on one side. Therefore, you need to get rid of the 6 and 5. So first you'll do: 5x + 6 = 4 -6 -6 5x=-2 Can you do the rest? :o)

OpenStudy (anonymous):

its algebra 1 and thank you guys so much!

OpenStudy (stamp):

@Butterfly16 doing it right, @meyjoux just giving away the answer

OpenStudy (anonymous):

by the method tranposition and adding/subtracting like terms :))

OpenStudy (anonymous):

oops forgot the 1 on 41 lol :o) 5x=35 :o)

OpenStudy (dmezzullo):

@meyjoux plz show all the work

OpenStudy (anonymous):

do you need it @BlondeGirl ?? or you get it already?

OpenStudy (anonymous):

transpose 6 to other side and change the sign 5X=41-6 add like terms 5X=35 divide both sides by 5 X=7